The quantity supplied of a commodity at a price of Birr 10 per unit is 500 units. Its price elasticity of supply is 0.5 A) Calculate the price at which its quantity supplied is 600 units. B) What quantity of this commodity will the seller supplied when the price rise at Birr 12 per unit? C) What quantity of this commodity will the sellers supply when price rises by Birr 4 per unit?


1
Expert's answer
2022-01-13T09:01:37-0500

Solution:

A.). Price elasticity of supply = "=\frac{\%\;change\; in\; quantity\; supplied}{\%\; change\; in\; price}"


"0.5 = (\frac{600 -500}{500 }) \div (\frac{x -10}{10 } )"

X = 14

The price at which its quantity supplied is 600 units = 14

 

B.). Price elasticity of supply = "=\frac{\%\;change\; in\; quantity\; supplied}{\%\; change\; in\; price}"


"0.5 = (\frac{x -500}{500 }) \div (\frac{12 -10}{10 } )"

X = 550

The quantity of this commodity that the seller will supply when the price rise at Birr 12 per unit = 550 units

 

C.). Price elasticity of supply = "=\frac{\%\;change\; in\; quantity\; supplied}{\%\; change\; in\; price}"

"0.5 = (\frac{x -500}{500 }) \div (\frac{40 -10}{10 } )"

X = 1,250

The quantity of this commodity that the seller will supply when the price rise by Birr 4 per unit = 1,250 units